OpenStudy (anonymous):

Help please.... A certain computer can perform 10^5 calculations per second. How many calculations can it perform in 10 seconds? explain....

OpenStudy (anonymous):

Like explain the steps please!!!!

OpenStudy (jango_in_dtown):

1sec=10^5 calculations so 10sec=10*10^5=10^6 calculations

OpenStudy (jango_in_dtown):

@krisnest

OpenStudy (anonymous):

jango_IN_DTOWN, dat wuz such a simple explanation, yet really concise, wow